# Unlocking massively parallel spectral proper orthogonal decompositions in the PySPOD package

### v1

@@ 5.2 Strong scalability
To perform the strong scalability analysis, we used a 2D version of the dataset presented in section 4.2, corresponding to the 10 hPa pressure level Hersbach et al. (2018). This dataset contains 727,584 time snapshots on a 1440$\times$721 grid and is 1,407 GB in size when stored on the disk in netCDF format. The data set contains a total of 755.4 billion data points in single-precision floating point, which amounts to 5,628 GB in memory once stored in double-precision floating-point for this study. As we scale from 256 up to 8,192 processes (64 to 2,048 nodes), this corresponds to between 21.99 GB and 0.68 GB, or 4,055 and 126 spatial points over time per process. This provides a broad range of scenarios. In all cases, all the temporal data is split into 1-year (8,760 snapshots) blocks.
@@
The FFT calculation (violet curve), the second most expensive part of the SPOD algorithm, exhibited surprising behavior. The speedup is measured at 50$\times$ when the number of processes is increased by 32$\times$ (from 256 to 8,192 processes). This phenomenon has been previously observed on the same supercomputer Dalcin et al. (2019) and can be explained by cache effects, frequency throttling, and possibly problem size-dependent NumPy optimizations.
@@
In terms of overall efficiency, we calculate it to be 74% when using 2,048 cores and 38% or lower once increasing to 4,096 cores and beyond. Even though the efficiency is only 15% when using 8,192 processes, we note that in this scenario, the total runtime is only around 127 seconds. The shortest runtime was achieved when using 4,096 processes and was around 98 seconds, 4 seconds faster than when using 2,048 processes. Based on those results and to use the computational resources responsibly, we suggest using at least 1 GB of data per process in practical applications.
@@ Acknowledgements
Marcin Rogowski, Lisandro Dalcin, and Matteo Parsani were supported by King Abdullah University of Science and Technology (KAUST). Brandon C. Y. Yeung and Oliver T. Schmidt gratefully acknowledge support from Office of Naval Research award N00014-23-1-2457, under the supervision of Dr. Steve Martens. Romit Maulik was supported by U.S. DOE ASCR Award Data-intensive Scientific Machine Learning: DE-FOA-2493. Gianmarco Mengaldo was supported by MOE Tier 2 grant 22-5191-A0001-0. The authors are thankful to the KAUST Supercomputing Laboratory for their computing resources. LES calculations were carried out on the “Onyx” Cray XC40/50 system in ERDC DSRC, using allocations provided by DoD HPCMP.

### v2

@@ 5.2 Strong scalability
To perform the strong scalability analysis, we used a 2D version of the dataset presented in section 4.2, corresponding to the 10 hPa pressure level [59]. This dataset contains 727,584 time snapshots on a 1440$\times$721 grid and is 1,407 GB in size when stored on the disk in netCDF format. The data set contains a total of 755.4 billion data points in single-precision floating point, which amounts to 5,628 GB in memory once stored in double-precision floating-point for this study. As we scale from 256 up to 8,192 processes (64 to 2,048 nodes), this corresponds to between 21.99 GB and 0.68 GB, or 4,055 and 126 spatial points over time per process. This provides a broad range of scenarios. In all cases, all the temporal data is split into 1-year (8,760 snapshots) blocks. This choice of block size is different to that of section 4.2 for the sake of using consistent configurations in both strong and weak scalability analyses.
@@
The FFT calculation (violet curve), the second most expensive part of the SPOD algorithm, exhibited surprising behavior. The speedup is measured at 50$\times$ when the number of processes is increased by 32$\times$ (from 256 to 8,192 processes). This phenomenon has been previously observed on the same supercomputer [60] and can be explained by cache effects, frequency throttling, and possibly problem size-dependent NumPy optimizations.
@@
In terms of overall relative efficiency [^1] The relative parallel efficiency is calculated as $E_{P}=(256\times T_{256})/(P\times T_{P})$, where $T_{P}$ is the wall-clock time corresponding to $P$ processes., we calculate it to be 74% when using 2,048 cores and 38% or lower once increasing to 4,096 cores and beyond. Even though the efficiency is only 15% when using 8,192 processes, we note that in this scenario, the total runtime is only around 127 seconds. The shortest runtime was achieved when using 4,096 processes and was around 98 seconds, 4 seconds faster than when using 2,048 processes. Based on the presented performance characteristics and emphasizing that I/O is the dominating cost, we suggest using as many processes as required to saturate the I/O bandwidth. Increasing the dedicated resources past the I/O saturation point will significantly reduce the parallel efficiency.
@@ Acknowledgements
Marcin Rogowski, Lisandro Dalcin, and Matteo Parsani acknowledge support from King Abdullah University of Science and Technology (KAUST) award BAS/1/1663-01-01. Brandon C. Y. Yeung and Oliver T. Schmidt gratefully acknowledge support from Office of Naval Research award N00014-23-1-2457, under the supervision of Dr. Steve Martens. Romit Maulik was supported by U.S. DOE ASCR Award Data-intensive Scientific Machine Learning: DE-FOA-2493. Gianmarco Mengaldo was supported by MOE Tier 2 grant 22-5191-A0001-0: “Prediction-to-Mitigation with Digital Twins of the Earth’s Weather”. The authors are thankful to the KAUST Supercomputing Laboratory for their computing resources. LES calculations were carried out on the “Onyx” Cray XC40/50 system in ERDC DSRC, using allocations provided by DoD HPCMP.